# PointSeg: A Training-Free Paradigm for 3D Scene Segmentation via Foundation Models

### v3

@@ Abstract
Recent success of vision foundation models have shown promising performance for the 2D perception tasks. However, it is difficult to train a 3D foundation network directly due to the limited dataset and it remains under explored whether existing foundation models can be lifted to 3D space seamlessly. In this paper, we present PointSeg, a novel training-free paradigm that leverages off-the-shelf vision foundation models to address 3D scene perception tasks. PointSeg can segment anything in 3D scene by acquiring accurate 3D prompts to align their corresponding pixels across frames. Concretely, we design a two-branch prompts learning structure to construct the 3D point-box prompts pairs, combining with the bidirectional matching strategy for accurate point and proposal prompts generation. Then, we perform the iterative post-refinement adaptively when cooperated with different vision foundation models. Moreover, we design a affinity-aware merging algorithm to improve the final ensemble masks. PointSeg demonstrates impressive segmentation performance across various datasets, all without training. Specifically, our approach significantly surpasses the state-of-the-art specialist training-free model by 14.1, 12.3, and 12.6 mAP on ScanNet, ScanNet++, and KITTI-360 datasets, respectively. On top of that, PointSeg can incorporate with various foundation models and even surpasses the specialist training-based methods by 3.4-5.4 mAP across various datasets, serving as an effective generalist model.
@@ 1 Introduction
To address these challenges, we present PointSeg, a novel perception framework that effectively incorporates different foundation models for tackling the 3D scene segmentation task without training. The key idea behind PointSeg is to learn accurate 3D point-box prompts pairs to enforce the off-the-shelf foundation models and fully unleash their potential in 3D scene segmentation tasks with three effective components. First, we construct a two-branch prompts learning structure to acquire the 3D point prompts and 3D box prompts respectively. The 3D point prompts are derived from localization abilities of PointLLM (Xu et al., 2024) to provide more explicit prompts in the form of points and 3D box prompts come from the 3D detectors (Shen et al., 2024; Wu et al., 2023b). Considering that the naive prompts could result in fragmented false-positive masks caused by matching outliers, we propose the bidirectional matching strategy for the generation of accurate point-box prompt pairs. Furthermore, when incorporated with different vision segmentation foundation models, such as SAM (Kirillov et al., 2023), our approach involves the iterative post-refinement to eliminate the coarse boundaries and isolated instances of background noise adaptively. Finally, with the primary aim of segmenting all points within the 3D scene, we employ the affinity-aware merging algorithm to capture pairwise similarity scores based on the 2D masks generated by the vision segmentation foundation models.
Comprehensive experiments on ScanNet (Dai et al., 2017), ScanNet++ (Yeshwanth et al., 2023), and KITTI-360 (Liao et al., 2022) demonstrate the superior generalization of the proposed PointSeg, surpassing previous specialist training-free model by 12.3-14.1 mAP and specialist training-based methods by 3.4-5.4 mAP across different datasets, all without training on domain-specific data. Remarkably, our zero-shot approach yields superior results in comparison to fully-supervised PointSAM (Zhou et al., 2024) trained on synthetic datasets, thereby emphasizing the effectiveness of PointSeg in the segmentation of intricate 3D scene. Moreover, we incorporate different foundation models, i.e., FastSAM (Zhao et al., 2023), MobileSAM (Zhang et al., 2023a), and EfficientSAM (Xiong et al., 2023), into our pipeline, and the performance gain shows that enhancements on 2D images can be seamlessly translated to improve 3D results. We summarize the contributions of our paper as follows:
@@ 2 Related Work
The emergence of the Segment Anything Model (SAM) (Kirillov et al., 2023) have triggered a revolution in the field of 2D segmentation. Having been trained on the extraordinary SA-1B dataset, SAM has garnered a vast amount of knowledge, equipping it to effectively segment unfamiliar images without additional training. Followed by SAM, several works (Zhang et al., 2023b; Zou et al., 2024; Xiong et al., 2023; Zhao et al., 2023; Zhang et al., 2023a; Liu et al., 2023) have attempted to accelerate or customize the original SAM from different aspects. Recognizing the exceptional capabilities of SAM, various recent research initiatives are working diligently to incorporate SAM into 3D scene segmentation task. Several (Yang et al., 2023b; Yin et al., 2024; Guo et al., 2023a) works attempt to segment each frame individually or constructs a graph based on the superpoints to lift the segmentation results to 3D space. However, these methods designate pixel prompts that are specific to each frame but do not synchronize across frames. This causes inconsistencies in segmentation across frames and produces substandard 3D segmentation results. Different from these 2D-to-3D lifting methods, SAMPro3D (Xu et al., 2023) attempts to locate 3D points in scenes as 3D prompts to align their projected pixel prompts across frames. Albeit effective, simply connecting 3D points to 2D space through projection is still too rough for complex scenes. In this paper, we propose a two-branch prompts learning structure towards accurate 3D prompts generation.
@@ 3 Method
As illustrated in Figure 2, we build a training-free 3D scene segmentation framework based on the off-the-shelf foundation models. Our PointSeg consists of three parts: 1) Bidirectional Matching based Prompts Generation (BMP) (Section 3.1). Given the reconstructed scene point cloud together with a set of posed RGB-D images , PointSeg first employs a two-branch prompts learning structure to acquire the 3D mask and box prompts, respectively. And the final point-box prompts pairs are obtained by the further bidirectional matching. These prompts serve as inputs to vision foundation models, such as SAM (Kirillov et al., 2023), after aligning with the pixels in 2D images. 2) Iterative 2D Mask Generation (Section 3.2). Then, we perform Iterative Post-refinement (IPR) to enable the generation of mask proposals adaptively. 3) Affinity-aware 3D Mask Refinement (Section 3.3). We calculate the affinity scores between the points generated in the point-box pairs and the mask proposals, followed by the Affinity-aware Merging (AM) (see Algorithm 2) to obtain the final 3D segmentation masks.
@@ 3.1 Bidirectional Matching based Prompts Generation
Apart from the point-level prompts in the 3D mask, we intend to inquire into how to generate dense region-level prompts to fully unleash the advantages of 3D prompts in the another branch. To enhance the ability to segment regions accurately in the subsequent segmenters, we exploit the localization abilities of 3D detector to provide more explicit prompts in the form of bounding boxes. The point cloud is taken as input by the frozen 3D detectors to generate the 3D bounding box for each category with corresponding proposal features , which can be represented as .
@@ 3.2 Iterative 2D Mask Generation
The vanilla segmentation model, such as SAM (Kirillov et al., 2023), accepts various inputs such as pixel coordinates, bounding boxes or masks and predict the segmentation area associated with each prompt. Hence, we feed the projected 2D point-box pairs calculated before into the foundation segmenters.
@@
Through the above operation, we can obtain 2D segmentation mask on all frames from the decoder, which however, might include rough edges and isolated background noises. For further refinement, we iteratively feed the mask back into the decoder for the adaptive post-processing. As illustrated in Algorithm 1, we first obtain the 2D mask by feeding the 2D point-box pairs into the SAM-based decoder. Then we prompt the decoder additionally with this mask along with these projected prompt pairs to obtain the next mask. And the initial value to record change is set to infinity. In each subsequent iteration, we calculate the change ratio between the two adjacent masks and compare it with our predefined threshold , which is set to 5 by default. We repeat this iterative process until the change value falls below the threshold adaptively. It is worth noting that we have also tried the method of fixed iteration steps, but this adaptive iteration method have proved to be more effective which is shown in Figure 4 and Table B.
@@ 3.3 Affinity-aware 3D Mask Refinement
Method Type ScanNet ScanNet++ KITTI-360 mAP AP50 AP25 mAP AP50 AP25 mAP AP50 AP25 With Training SAM-graph (Guo et al., 2023a) specialist model 15.1 33.3 59.1 12.9 25.3 43.6 14.7 28.0 43.2 Mask3D (Schult et al., 2023) specialist model 26.9 44.4 57.5 8.8 15.0 22.3 0.1 0.4 4.2 OpenDAS (Yilmaz et al., 2024) specialist model 28.3 49.6 66.2 19.2 35.5 52.6 20.1 32.4 52.2 PointSAM (Zhou et al., 2024) specialist model 32.9 56.4 72.5 25.8 38.0 59.3 25.1 38.4 56.2 Training-free SAM3D (Yang et al., 2023b) specialist model 13.7 29.7 54.5 8.3 17.5 33.7 6.3 16.0 35.6 SAI3D (Yin et al., 2024) specialist model 18.8 42.5 62.3 17.1 31.1 49.5 16.5 30.2 48.6 SAMPro3D (Xu et al., 2023) specialist model 22.2 45.6 65.7 18.9 33.7 51.6 17.3 31.1 49.6 PointSeg(ours) generalist model 36.3 60.2 79.3 31.2 46.5 64.8 29.9 44.5 63.3
Method ScanNet ScanNet++ KITTI-360 mAP AP50 AP25 mAP AP50 AP25 mAP AP50 AP25 + SAM (Kirillov et al., 2023) 36.3 60.2 79.3 31.2 46.5 64.8 29.9 44.5 63.3 + MobileSAM (Zhang et al., 2023a) 26.2 49.8 68.3 19.6 36.4 55.2 20.6 34.6 53.3 + FastSAM (Zhao et al., 2023) 26.9 50.8 69.1 20.5 37.7 56.5 21.2 35.8 54.5 + EfficientSAM (Xiong et al., 2023) 33.5 57.2 75.8 27.8 43.6 62.4 27.5 41.7 60.6
@@ 4.2 Main Results
We compare the segmentation results on ScanNet, ScanNet++, and KITTI-360 datasets, covering both indoor and outdoor scenes. As shown in Table 1, comparing to previous training-free methods, our PointSeg obtains 14.1 mAP, 14.6 AP50, and 13.6 AP25 performance gains on ScanNet. On the more challenging indoor dataset ScanNet++, our method still obtains 12.3 mAP, 12.8 AP50, and 13.2 AP25 improvements. Furthermore, when evaluating the performance of our PointSeg on outdoor KITTI-360, our method still surpasses corresponding zero-shot method by 12.6 mAP, 13.4 AP50, and 13.7 AP25, respectively. In this regard, our PointSeg demonstrates superior generalization ability to complex 3D scenarios.
Notably, when compared to previous training-based methods, PointSeg outperforms PointSAM (Zhou et al., 2024) by 3.4-5.4, 3.8-6.5, 5.5-7.1 in terms of mAP, AP50, and AP25 across various datasets. This further demonstrates the robustness and effectiveness of our approach in 3D segmentation task.
@@ 4.3 Ablation Study
In this section, we conduct extensive ablation studies on ScanNet to show the effectiveness of each component and design. More ablation studies are presented in Appendix B.
@@
Different Foundation Models. i) Apart the basic segmentation foundation model SAM (Kirillov et al., 2023), we also integrate different segmentation foundation model, i.e., MobileSAM (Zhang et al., 2023a), FastSAM (Zhao et al., 2023), and EfficientSAM (Xiong et al., 2023), into our framework. As shown in Table 2, PointSeg demonstrates consistent performance improvements among different datasets, where the original SAM-based result performs best. This is consistent with the relative results of these methods in 2D segmentation. ii) In Table 4.1, we change the model for localization of point prompts. The results demonstrate that more accurate points for 3D prompts can indeed contribute to the performance gain. ii) In Table 4.1, the improvement of different 3D detectors will also bring about the improvement of the performance of our method. These results suggest that our framework can serve as a foundational structure, capable of integrating a variety of fundamental models. And the enhancements observed in these models can be smoothly translated to 3D space, thereby augmenting the overall performance.
Main Components. Further, we explore the effects of bidirectional matching based prompts generation (BMP), iterative post-refinement (IPR) and affinity-aware merging (AM). We illustrate the importance of different components by removing some parts and keeping all the others unchanged. The baseline setting is to use points and boxes independently as 3D prompts. And the masks from the 2D segmentation model decoder are used to merge the final 3D masks, without any post-refinement. The merging strategy is same as the adjacent frame merging in (Yang et al., 2023b). The results of components ablations are shown in Table 5. We observe that using the iterative post-refinement strategy leads to a noticeable increase in performance, which demonstrates the necessity of refining the initial 2D masks. The performance degradation caused by the absence of bidirectional matching proves that the constraints between the point and box prompts can indeed help to generate the accurate point-box pairs. And the performance drop without the affinity-aware merging shows that the affinity score is indeed useful to link the point and the masks.

### v4

@@ Abstract
Recent success of vision foundation models have shown promising performance for the 2D perception tasks. However, it is difficult to train a 3D foundation network directly due to the limited dataset and it remains under explored whether existing foundation models can be lifted to 3D space seamlessly. In this paper, we present PointSeg, a novel training-free paradigm that leverages off-the-shelf vision foundation models to address 3D scene perception tasks. PointSeg can segment anything in 3D scene by acquiring accurate 3D prompts to align their corresponding pixels across frames. Concretely, we design a two-branch prompts learning structure to construct the 3D point-box prompts pairs, combining with the bidirectional matching strategy for accurate point and proposal prompts generation. Then, we perform the iterative post-refinement adaptively when cooperated with different vision foundation models. Moreover, we design a affinity-aware merging algorithm to improve the final ensemble masks. PointSeg demonstrates impressive segmentation performance across various datasets, all without training. Specifically, our approach significantly surpasses the state-of-the-art specialist training-free model by 16.3, 14.9, and 15 mAP on ScanNet, ScanNet++, and KITTI-360 datasets, respectively. On top of that, PointSeg can incorporate with various foundation models and even surpasses the specialist training-based methods by 5.6-8 mAP across various datasets, serving as an effective generalist model.
@@ 1 Introduction
To address these challenges, we present PointSeg, a novel perception framework that effectively incorporates different foundation models for tackling the 3D scene segmentation task without training. The key idea behind PointSeg is to learn accurate 3D point-box prompts pairs to enforce the off-the-shelf foundation models and fully unleash their potential in 3D scene segmentation tasks with three effective components. First, we construct a two-branch prompts learning structure to acquire the 3D point prompts and 3D box prompts respectively. The 3D point prompts are derived from localization abilities of PointLLM (Xu et al., 2024) to provide more explicit prompts in the form of points and 3D box prompts come from the 3D detectors (Shen et al., 2024; Wu et al., 2023b). Considering that the naive prompts could result in fragmented false-positive masks caused by matching outliers, we propose the bidirectional matching strategy for the generation of accurate point-box prompt pairs. Furthermore, when incorporated with different 2D vision segmentation foundation models, such as SAM 2 (Ravi et al., 2024), our approach involves the iterative post-refinement to eliminate the coarse boundaries and isolated instances of background noise adaptively. Finally, with the primary aim of segmenting all points within the 3D scene, we employ the affinity-aware merging algorithm to capture pairwise similarity scores based on the 2D masks generated by the vision segmentation foundation models.
Comprehensive experiments on ScanNet (Dai et al., 2017), ScanNet++ (Yeshwanth et al., 2023), and KITTI-360 (Liao et al., 2022) demonstrate the superior generalization of the proposed PointSeg, surpassing previous specialist training-free model by 14.9-16.3 mAP and specialist training-based methods by 5.6-8 mAP across different datasets, all without training on domain-specific data. Remarkably, our zero-shot approach yields superior results in comparison to fully-supervised PointSAM (Zhou et al., 2024) trained on synthetic datasets, thereby emphasizing the effectiveness of PointSeg in the segmentation of intricate 3D scene. Moreover, we incorporate different foundation models, i.e., SAM 2 (Ravi et al., 2024), SAM (Kirillov et al., 2023), FastSAM (Zhao et al., 2023), MobileSAM (Zhang et al., 2023a), and EfficientSAM (Xiong et al., 2023), into our pipeline, and the performance gain shows that enhancements on 2D images can be seamlessly translated to improve 3D results. We summarize the contributions of our paper as follows:
@@ 2 Related Work
The emergence of the Segment Anything Model (Kirillov et al., 2023; Ravi et al., 2024) have triggered a revolution in the field of 2D segmentation. Having been trained on the extraordinary SA-1B dataset, SAM has garnered a vast amount of knowledge, equipping it to effectively segment unfamiliar images without additional training. Followed by SAM, several works (Zhang et al., 2023b; Zou et al., 2024; Xiong et al., 2023; Zhao et al., 2023; Zhang et al., 2023a; Liu et al., 2023) have attempted to accelerate or customize the original SAM from different aspects. Recognizing the exceptional capabilities of SAM, various recent research initiatives are working diligently to incorporate SAM into 3D scene segmentation task. Several works (Yang et al., 2023b; Yin et al., 2024; Guo et al., 2023a) attempt to segment each frame individually or constructs a graph based on the superpoints to lift the segmentation results to 3D space. However, these methods designate pixel prompts that are specific to each frame but do not synchronize across frames. This causes inconsistencies in segmentation across frames and produces substandard 3D segmentation results. Different from these 2D-to-3D lifting methods, SAMPro3D (Xu et al., 2023) attempts to locate 3D points in scenes as 3D prompts to align their projected pixel prompts across frames. Albeit effective, simply connecting 3D points to 2D space through projection is still too rough for complex scenes. In this paper, we propose a two-branch prompts learning structure towards accurate 3D prompts generation.
@@ 3 Method
As illustrated in Figure 2, we build a training-free 3D scene segmentation framework based on the off-the-shelf foundation models. Our PointSeg consists of three parts: 1) Bidirectional Matching based Prompts Generation (BMP) (Section 3.1). Given the reconstructed scene point cloud together with a set of posed RGB-D images , PointSeg first employs a two-branch prompts learning structure to acquire the 3D mask and box prompts, respectively. And the final point-box prompts pairs are obtained by the further bidirectional matching. These prompts serve as inputs to 2D vision foundation models, such as SAM 2 (Ravi et al., 2024), after aligning with the pixels in 2D images. 2) Iterative 2D Mask Generation (Section 3.2). Then, we perform Iterative Post-refinement (IPR) to enable the generation of mask proposals adaptively. 3) Affinity-aware 3D Mask Refinement (Section 3.3). We calculate the affinity scores between the points generated in the point-box pairs and the mask proposals, followed by the Affinity-aware Merging (AM) (see Algorithm 2) to obtain the final 3D segmentation masks.
@@ 3.1 Bidirectional Matching based Prompts Generation
Apart from the point-level prompts in the 3D mask, we intend to inquire into how to generate dense region-level prompts to fully unleash the advantages of 3D prompts in the another branch. To enhance the ability to segment regions accurately in the subsequent segmenters, we exploit the localization abilities of 3D detector to provide more explicit prompts in the form of bounding boxes. The point cloud is taken as input by the frozen 3D detectors to generate the 3D bounding box for each category with corresponding proposal features , which can be represented as
@@ 3.2 Iterative 2D Mask Generation
The vanilla segmentation model, such as SAM 2 (Ravi et al., 2024), accepts various inputs such as pixel coordinates, bounding boxes or masks and predict the segmentation area associated with each prompt. Hence, we feed the projected 2D point-box pairs calculated before into the foundation segmenters.
@@
Through the above operation, we can obtain 2D segmentation mask on all frames from the decoder, which however, might include rough edges and isolated background noises. For further refinement, we iteratively feed the mask back into the decoder for the adaptive post-processing. As illustrated in Algorithm 1, we first obtain the 2D mask by feeding the 2D point-box pairs into the SAM-based decoder. Then we prompt the decoder additionally with this mask along with these projected prompt pairs to obtain the next mask. And the initial value to record change is set to infinity. In each subsequent iteration, we calculate the change ratio between the two adjacent masks and compare it with our predefined threshold , which is set to 5 by default. We repeat this iterative process until the change value falls below the threshold adaptively. It is worth noting that we have also tried the method of fixed iteration steps, but this adaptive iteration method have proved to be more effective which is shown in Figure 4 and Table 4.3.
@@ 3.3 Affinity-aware 3D Mask Refinement
Method Type ScanNet ScanNet++ KITTI-360 mAP AP50 AP25 mAP AP50 AP25 mAP AP50 AP25 With Training SAM-graph (Guo et al., 2023a) specialist model 15.1 33.3 59.1 12.9 25.3 43.6 14.7 28.0 43.2 Mask3D (Schult et al., 2023) specialist model 26.9 44.4 57.5 8.8 15.0 22.3 0.1 0.4 4.2 OpenDAS (Yilmaz et al., 2024) specialist model 28.3 49.6 66.2 19.2 35.5 52.6 20.1 32.4 52.2 PointSAM (Zhou et al., 2024) specialist model 32.9 56.4 72.5 25.8 38.0 59.3 25.1 38.4 56.2 Training-free SAM3D (Yang et al., 2023b) specialist model 13.7 29.7 54.5 8.3 17.5 33.7 6.3 16.0 35.6 SAI3D (Yin et al., 2024) specialist model 18.8 42.5 62.3 17.1 31.1 49.5 16.5 30.2 48.6 SAMPro3D (Xu et al., 2023) specialist model 22.2 45.6 65.7 18.9 33.7 51.6 17.3 31.1 49.6 PointSeg (Ours) generalist model 38.5 63.6 82.1 33.8 49.1 67.2 32.3 47.2 66.5
Method ScanNet ScanNet++ KITTI-360 mAP AP50 AP25 mAP AP50 AP25 mAP AP50 AP25 + SAM 2 (Ravi et al., 2024) 38.5 63.6 82.1 33.8 49.1 67.2 32.3 47.2 66.5 + SAM (Kirillov et al., 2023) 36.3 60.2 79.3 31.2 46.5 64.8 29.9 44.5 63.3 + MobileSAM (Zhang et al., 2023a) 26.2 49.8 68.3 19.6 36.4 55.2 20.6 34.6 53.3 + FastSAM (Zhao et al., 2023) 26.9 50.8 69.1 20.5 37.7 56.5 21.2 35.8 54.5 + EfficientSAM (Xiong et al., 2023) 33.5 57.2 75.8 27.8 43.6 62.4 27.5 41.7 60.6
@@ 4.2 Main Results
We compare the segmentation results on ScanNet, ScanNet++, and KITTI-360 datasets, covering both indoor and outdoor scenes. As shown in Table 1, comparing to previous training-free methods, our PointSeg obtains 16.3 mAP, 18 AP50, and 16.4 AP25 performance gains on ScanNet. On the more challenging indoor dataset ScanNet++, our method still obtains 14.9 mAP, 15.4 AP50, and 15.6 AP25 improvements. Furthermore, when evaluating the performance of our PointSeg on outdoor KITTI-360, our method still surpasses corresponding zero-shot method by 15 mAP, 16.1 AP50, and 16.9 AP25, respectively. In this regard, our PointSeg demonstrates superior generalization ability to complex 3D scenarios.
Notably, when compared to previous training-based methods, PointSeg outperforms PointSAM (Zhou et al., 2024) by 5.6-8, 7.2-11.1, 7.9-10.3 in terms of mAP, AP50, and AP25 across various datasets. This further demonstrates the robustness and effectiveness of our approach in 3D segmentation task.
@@ 4.3 Ablation Study
In this section, we conduct extensive ablation studies on ScanNet to show the effectiveness of each component and design. Unless otherwise specified, SAM (Kirillov et al., 2023) is used as 2D segmentation foundation model for ablation studies by default.
Different Foundation Models. i) Apart the basic segmentation foundation model SAM 2 (Ravi et al., 2024), we also integrate different segmentation foundation model, i.e., SAM (Kirillov et al., 2023), MobileSAM (Zhang et al., 2023a), FastSAM (Zhao et al., 2023), and EfficientSAM (Xiong et al., 2023), into our framework. As shown in Table 2, PointSeg demonstrates consistent performance improvements among different datasets, where the original SAM 2-based result performs best. This is consistent with the relative results of these methods in 2D segmentation. ii) In Table 4.1, we change the model for localization of point prompts. The results demonstrate that more accurate points for 3D prompts can indeed contribute to the performance gain. iii) In Table 4.1, the improvement of different 3D detectors will also bring about the improvement of the performance of our method. These results suggest that our framework can serve as a foundational structure, capable of integrating a variety of fundamental models. And the enhancements observed in these models can be smoothly translated to 3D space, thereby augmenting the overall performance.
Main Components. Further, we explore the effects of bidirectional matching based prompts generation (BMP), iterative post-refinement (IPR) and affinity-aware merging (AM). We illustrate the importance of different components by removing some parts and keeping all the others unchanged. The baseline setting is to use points and boxes independently as 3D prompts. And the masks from the 2D segmentation model decoder are used to merge the final 3D masks, without any post-refinement. The merging strategy is same as the adjacent frame merging in (Yang et al., 2023b). The results of components ablations are shown in Table 4.3. We observe that using the iterative post-refinement strategy leads to a noticeable increase in performance, which demonstrates the necessity of refining the initial 2D masks. The performance degradation caused by the absence of bidirectional matching proves that the constraints between the point and box prompts can indeed help to generate the accurate point-box pairs. And the performance drop without the affinity-aware merging shows that the affinity score is indeed useful to link the point and the masks.
@@
Iterative Post-refinement. As mentioned before, when performing the post-refinement during the initial 2D mask generation, we have tried different strategies of fixed numbers method and adaptive iteration method. As shown in Table 4.3, the six rows in the middle represent the method to use a fixed number of iterations and the last row is the adaptive iteration. We can observe that as the number of iterations increases, the corresponding AP value also becomes higher compared to no iteration, which shows that the obtained mask is also more accurate. In the method with a fixed number of iterations, the results reach the highest in the third iteration, but are still lower than those in the adaptive iteration method. This largely illustrates the effectiveness of the iterative post-processing method in generating more accurate masks, and also reveals that the adaptive iteration is more beneficial.
Different Matching Strategy. To validate the effect of different matching method, we explore the effects of the forward matching and the reverse matching of the proposed bidirectional matching, as shown in Table 4.3. Without the guidance from the respective point and box, the naive prompts contain many invalid points and regions, which provide negative prompts for the following segmentation models. Our bidirectional matching improves the performance of forward and reverse matching by 5.4 and 3.7 , which show the effectiveness of the proposed bidirectional matching strategy.
module bidirectional matching iterative refinement affinity-aware merging FPS 1.53 1.05 1.96
Affinity-aware Merging. In the final mask merging stage, we have also tried other merging algorithm. As shown in Table 4.3, we compare our affinity-aware merging (AM) with (a) bidirectional merging (BM) from (Yang et al., 2023b), (b) pure merging (PM) without affinity scores, which is simplified from our approach and (c) prompt ID based merging (IDM) from (Xu et al., 2023). With other inferior merging method, the performance drops dramatically which shows the superiority of our proposed affinity-aware merging algorithm in solving the mask merging problems in 3D scene.
3D Point Models. As shown in Table 4.3, the missing of PointLLM/3D detector causes the performance drop and only the matching of combination performs best.
Mask Change Ratio. In the iterative post-refinement module, we set a mask change ratio threshold as a condition for stopping the iteration. Here, We show the effect of different ratios on the results. As shown in Table 10, the overall results perform best when mask change ratio is set as 5.
Inference Speed. As shown in Table 11, we test the running efficiency on NVIDIA V100 GPU and the detailed FPS of each module. The FLOPs of our method is 4.2G.